OpenStudy (anonymous):

WHAT IS THE VERTEX OF THE PARABOLA DESCRIBED BY Y=X^2-12X+3

OpenStudy (jdoe0001):

so yo have the so-called standard equation form you'd need to change it to the VERTEX form http://tccl.rit.albany.edu/knilt/images/1/12/Standard-vertex-forms.png

OpenStudy (jdoe0001):

so, do you know what a perfect square trinomial is?

OpenStudy (anonymous):

NO MY TEACHER DIDNT SAY ANY THING ABOUT THAT

OpenStudy (johnweldon1993):

Well the 'x' coordinate can be found by using the equation \[-\frac{ b }{ 2a }\] you have an equation in standard form ax² + bx + c x² -12x + 3 so a = 1 b = -12 c = 3 Plug them into that equation \[\frac{ 12 }{ 2 }\] Now you can see the vertex 'x' coordinate = 6 To find the 'y' coordinate....plug (6) in for 'x' of your original equation Y = x² - 12x + 3 Y = 6² - 12(6) + 3 Y = ...?

OpenStudy (anonymous):

SO MY ANSWER IS (6,-33)

OpenStudy (johnweldon1993):

That would be correct...those are the coordinates of the vertex

OpenStudy (johnweldon1993):

And I hope you see how I did it and it is clear :)

OpenStudy (anonymous):

THANK YOU I WQROTE DOWN THE WHOLE THING IN MY NOTES TO HAVE FOR MY OTHER PROBLEMS

OpenStudy (johnweldon1993):

Perfect! :) and you're welcome!